Pierre bikes 7 km/h faster than Sal. They start at the same place and travel in opposite directions for 2 hours. They are then 86 km apart. How fast is Pierre traveling?
Answer by richwmiller(17219)

We are asked for Pierre's speed so we solve for Pierre's speed p
p=s+7
s=p-7
2*p+2*(p-7)=86